# Pursuing Overall Welfare in Federated Learning through Sequential Decision Making

### v1

@@ 1 Introduction
Since each client updates the copy of a global model with its own data, variability across clients’ data distributions causes many problems (Kairouz et al., 2021; Li et al., 2020b). The client-level fairness (Chen et al., 2023) is one of the main problems affected by such a statistical heterogeneity (Kairouz et al., 2021; Li et al., 2020b; Mohri et al., 2019; Li et al., 2019). Although the performance of a global model is high in average, some clients may be more benefited than others, resulting in violation of the client-level fairness. In this situation, there inevitably exists a group of clients who cannot utilize the trained global model due to its poor performance. This is a critical problem in practice since the underperformed groups may lose motivation to participate in the FL system. To remedy this problem, previous works (Mohri et al., 2019; Li et al., 2019, 2020a; Hu et al., 2022; Zhang et al., 2022) proposed to modify the static aggregation scheme into an adaptive aggregation strategy, according to given local signals (e.g., losses or gradients). In detail, the server re-weights local updates by assigning larger mixing coefficients to higher local losses.
When updating the mixing coefficients, however, only a few bits are provided to the server, compared to the update of a model parameter. For example, suppose that there exist $K$ clients in the FL system, each of which has $N$ local samples. When all clients participate in each round, $KN$ samples are used effectively for updating a new global model $\boldsymbol{\theta}$. On the contrary, only $K$ bits (e.g., local losses: $F_{1}(\boldsymbol{\theta}),...,F_{K}(\boldsymbol{\theta})$) are provided to the server for an update of mixing coefficients. This is aggravated in cases where $K$ is too large, thus client sampling is inevitably required. In this case, the server is provided with far less than $K$ signals, which hinders faithful update of the mixing coefficients.
@@ 2 Related Works
The OCO framework is designed for making sequential decisions with the best utilities, having solid theoretical backgrounds. It aims to minimize the cumulative mistakes of a decision maker (e.g., central sever), given a response of the environment (e.g., losses from clients) for finite rounds $t\in[T]$. The cumulative mistakes of the learner are usually denoted as the cumulative regret (see (5)), and the learner can achieve sublinear regret in finite rounds using well-designed OCO algorithms (Shalev-Shwartz et al., 2012; McMahan, 2017; Orabona, 2019). In designing an OCO algorithm, two main frameworks are mainly considered: Online Mirror Descent (OMD) (Nemirovskij & Yudin, 1983; Warmuth et al., 1997; Beck & Teboulle, 2003) and Follow-The-Regularized-Leader (FTRL) (Abernethy et al., 2009; Hazan & Kale, 2010; Agarwal & Hazan, 2005; Shalev-Shwartz & Singer, 2006). One of popular instantiations of both frameworks is the Online Portfolio Selection (OPS) algorithm, of which decision space is restricted to a probability simplex. The universal portfolio algorithm is the first that yields an optimal theoretical regret, $\mathcal{O}(K\log{T})$) despite its heavy computation ($\mathcal{O}(K^{4}T^{14})$) (Cover, 1991), the Online Gradient Descent (Zinkevich, 2003) and the Exponentiated Gradient (EG) (Helmbold et al., 1998) show slightly worse regrets (both are $\mathcal{O}(\sqrt{T})$), but can be executed in linear runtime in practice ($\mathcal{O}(K)$). Plus, the Online Newton Step (ONS) (Agarwal et al., 2006; Hazan et al., 2007) presents logarithmic regret with quadratic runtime in $K$. Since these OPS algorithms are proven to perform well when the decision is a probability vector, we adopt them for finding adaptive mixing coefficients to achieve performance fairness in FL. To the best of our knowledge, we are the first to consider fair FL algorithms under the OCO framework.
@@ 4.1 Improved Design for Better Decision Making
Additionally, when equipped with the negative logarithmic growth as a decision loss (i.e., (4)), the problem c can be addressed due to its strict convexity and Lipscthiz continuity. (See Lemma 4.1) Note that when the loss function is convex, we can run the FTRL with a linearized loss (i.e., ~$\tilde{\ell}^{(t)}\left(\boldsymbol{p}\right)=\left\langle\boldsymbol{p},\boldsymbol{g}^{(t)}\right\rangle$ where $\boldsymbol{g}^{(t)}=\nabla\ell^{(t)}\left(\boldsymbol{p}^{(t)}\right)$). This is useful in that a closed-form update can be obtained thanks to the properties of the Fenchel conjugate (see Remark 4.5).
@@ 4.2.1 AAggFF-S: Algorithm for the Cross-Silo Federated Learning
In the cross-silo setting, it is typically assumed that all $K$ clients participate in $T$ rounds, since there are a moderately small number of clients in the FL system. Therefore, the server’s stateful decision making is beneficial for enhancing overall welfare across federation rounds. This is also favorable since existing OPS algorithms can be readily adopted.
@@ 4.2.2 AAggFF-D: Algorithm for the Cross-Device Federated Learning
The ONS has regret proportional to $K$ and runs in $\mathcal{O}\left(K^{2}+K^{3}\right)$ [^1] A generalized projection required for the Online Newton Step that can be solved in $\tilde{\mathcal{O}}\left(K^{3}\right)$ (Agarwal et al., 2006; Hazan et al., 2007). per round, which is nearly impossible to be adopted for the cross-device FL setting due to large $K$, even though the logarithmic regret is guaranteed in $T$. Instead, we can exploit the variant of EG adapted to FTRL (Orabona, 2019), which can be run in $\mathcal{O}\left(K\right)$ time per round.
@@
The large number of clients coerces the FL system to introduce the client sampling scheme in each round. Therefore, the decision maker (i.e., the central server) cannot always observe all entries of a response vector per round. This is problematic in terms of OCO, since OCO algorithms assume that they can acquire intact response vector for every round $t\in[T]$. Instead, when the client sampling is introduced, the learner can only observe entries of sampled client indices in the round $t$, denoted as $S^{(t)}$.
@@
Denote $C=P\left(i\in S^{(t)}\right)$ as a client sampling probability of a cross-device FL system for every round $t\in[T]$. The DR estimator $\breve{\boldsymbol{r}}^{(t)}$, of which element is defined in (10) is an unbiased estimator of given partially observed response vector $\boldsymbol{r}^{(t)}$. i.e., $\mathbb{E}\left[\breve{\boldsymbol{r}}^{(t)}\right]=\boldsymbol{r}^{(t)}$.
@@
Further denote $\breve{\boldsymbol{g}}$ as a gradient estimate from (11) using the DR estimator of a response vector according to (10),k at an arbitrary reference $\boldsymbol{r}_{0}$. Then, $\breve{\boldsymbol{g}}$ is an unbiased estimator of the linearized gradient of a decision loss at $\boldsymbol{r}_{0}$, which is close to the true gradient: $\mathbb{E}\left[\breve{\boldsymbol{g}}\right]=\tilde{\boldsymbol{g}}\approx\boldsymbol{g}$.
@@ 6 Experimental Results
We conduct experiments on datasets mirroring realistic scenarios in FL systems: multiple modalities (vision, text, speech, and tabular form) and natural data partitioning. We briefly summarize FL settings of each dataset in Table 4. For the cross-silo setting, we used Berka tabular dataset (Berka, 1999), MQP clinical text dataset (McCreery et al., 2020), and ISIC oncological image dataset (Codella et al., 2018) (also a part of FLamby benchmark (Ogier du Terrail et al., 2022)). For the cross-device setting, we used CelebA vision dataset (Liu et al., 2015), Reddit text dataset (both are parts of LEAF benchmark (Caldas et al., 2019)) and SpeechCommands audio dataset (Warden, 2018).
@@ 7 Limitations and Future Works
Our work suggests interesting future directions for a better FL system, which may also be a limitation of the current work. First, we can exploit side information (e.g., parameters of local updates) to not preserve all clients’ mixing coefficients, and filter out malicious signals for robustness. For example, the former can be realized by adopting other decision making schemes such as contextual Bayesian optimization (Char et al., 2019), and the latter can be addressed by clustered FL (Ghosh et al., 2020; Sattler et al., 2020) for a group-wise estimation of mixing coefficients. Both directions are promising and may improve the practicality of FL systems. Furthermore, the FTRL objective can be replaced by the Follow-The-Perturbed-Leader (FTPL) (Kalai & Vempala, 2005), of which random perturbation in decision making process can be directly linked to the differential privacy (DP (Dwork, 2006)) guarantee (McMillan, 2018), which is frequently considered for the cross-silo setting. Last but not least, further convergence analysis is required w.r.t. the parameter perspective along with mixing coefficients, e.g., resorting to bi-level optimization framework.
@@ Acknowledgements
Federated Learning (FL) poses a potential risk of skewed performance distributions toward partial sub-groups of clients. We believe that our proposed method can mitigate such an unfair situation in the FL system, thereby enhancing the trustworthiness of the FL system and social welfare. A potential concern is the situation where malicious clientd deliberately inflates their local signals to cause global updates to be biased towards themselves. To prevent such a catastrophic situation, we aim to improve the current work by exploiting information other than local losses, as mentioned in the Limitations and future work section. Authors are not aware of other critical ethical/social implications otherwise, but are open to discussing them if exist.
@@ C.1 Datasets
For the main experiments in Table 2 and Table 3, we used 3 datasets for the cross-silo setting (Berka (Berka, 1999), MQP (McCreery et al., 2020), and ISIC (Codella et al., 2018; Ogier du Terrail et al., 2022)), and 3 other datasets for the cross-device setting (CelebA, Reddit (Caldas et al., 2019), and SpeechCommands (Warden, 2018)). In this section, we describe details of each dataset about its task, metrics, and client partitioning. The statistics of all federated benchmarks are summarized in Table A2. Note that Avg. and Std. in the table refer to the average and a standard deviation of a sample size of each client in the FL system
@@ 1. $\bullet$
    CelebA is a vision dataset containing the facial images of celebrities (Liu et al., 2015). It is curated for federated setting in LEAF benchmark (Caldas et al., 2019), and is targeted for the binary classification task (presence of smile). We follow the processing of (Caldas et al., 2019), thereby each client corresponds to each celebrity, having 9,343 total clients in the FL system. Finally, we used top-1 accuracy for the evaluation metric following (Caldas et al., 2019).
@@ D.1 Pseudocode for ClientUpdate
  Return: $F_{i}\left(\boldsymbol{\theta}\right)$, $\boldsymbol{\theta}-\boldsymbol{\theta}^{(t+1)}_{i}$.

### v2

@@ 1 Introduction
Since each client updates the copy of a global model with its own data, variability across clients’ data distributions causes many problems (Kairouz et al., 2021; Li et al., 2020b). The client-level fairness (Chen et al., 2023) is one of the main problems affected by such a statistical heterogeneity (Kairouz et al., 2021; Li et al., 2020b; Mohri et al., 2019; Li et al., 2019). Although the performance of a global model is high in average, some clients may be more benefited than others, resulting in violation of the client-level fairness. In this situation, there inevitably exists a group of clients who cannot utilize the trained global model due to its poor performance. This is a critical problem in practice since the underperformed groups may lose motivation to participate in the federated system. To remedy this problem, previous works (Mohri et al., 2019; Li et al., 2019, 2020a; Hu et al., 2022; Zhang et al., 2022) proposed to modify the static aggregation scheme into an adaptive aggregation strategy, according to given local signals (e.g., losses or gradients). In detail, the server re-weights local updates by assigning larger mixing coefficients to higher local losses.
When updating the mixing coefficients, however, only a few bits are provided to the server, compared to the update of a model parameter. For example, suppose that there exist $K$ clients in the federated system, each of which has $N$ local samples. When all clients participate in each round, $KN$ samples are used effectively for updating a new global model $\boldsymbol{\theta}$. On the contrary, only $K$ bits (e.g., local losses: $F_{1}(\boldsymbol{\theta}),...,F_{K}(\boldsymbol{\theta})$) are provided to the server for an update of mixing coefficients. This is aggravated in cases where $K$ is too large, thus client sampling is inevitably required. In this case, the server is provided with far less than $K$ signals, which hinders faithful update of the mixing coefficients.
@@ 2 Related Works
The OCO framework is designed for making sequential decisions with the best utilities, having solid theoretical backgrounds. It aims to minimize the cumulative mistakes of a decision maker (e.g., central sever), given a response of the environment (e.g., losses from clients) for finite rounds $t\in[T]$. The cumulative mistakes of the learner are usually denoted as the cumulative regret (see (5)), and the learner can achieve sublinear regret in finite rounds using well-designed OCO algorithms (Shalev-Shwartz et al., 2012; McMahan, 2017; Orabona, 2019). In designing an OCO algorithm, two main frameworks are mainly considered: Online Mirror Descent (OMD) (Nemirovskij & Yudin, 1983; Warmuth et al., 1997; Beck & Teboulle, 2003) and Follow-The-Regularized-Leader (FTRL) (Abernethy et al., 2009; Hazan & Kale, 2010; Agarwal & Hazan, 2005; Shalev-Shwartz & Singer, 2006). One of popular instantiations of both frameworks is the Online Portfolio Selection (OPS) algorithm, of which decision space is restricted to a probability simplex. The universal portfolio algorithm is the first that yields an optimal theoretical regret, $\mathcal{O}(K\log{T})$) despite its heavy computation ($\mathcal{O}(K^{4}T^{14})$) (Cover, 1991), the Online Gradient Descent (Zinkevich, 2003) and the Exponentiated Gradient (EG) (Helmbold et al., 1998) show slightly worse regrets (both are $\mathcal{O}(\sqrt{T})$), but can be executed in linear runtime in practice ($\mathcal{O}(K)$). Plus, the Online Newton Step (ONS) (Agarwal et al., 2006; Hazan et al., 2007) presents logarithmic regret with quadratic runtime in $K$. Since these OPS algorithms are proven to perform well when the decision is a probability vector, we adopt them for finding adaptive mixing coefficients to achieve performance fairness in FL. Apart from OPS, we refer interested readers to fair resource allocation, which allocates limited resources to users in a network system (Kelly et al., 1998; Altman et al., 2008; Joe-Wong et al., 2013) especially in dynamic manner (Si Salem et al., 2022; Banerjee et al., 2023; Talebi & Proutiere, 2018). To the best of our knowledge, we are the first to consider fair FL algorithms under the OCO framework.
@@ 3.3 Sequential Probability Assignment
It is also intuitive that the negative logarithmic growth is a soft approximation of maximum operator (i.e., log-sum-exp trick, or soft-max). Thus, the minimization of (4) w.r.t. $\boldsymbol{p}$ is equal to the maximization of the soft-max approximation of $\sum_{i=1}^{K}p_{i}r_{i}^{(t)}\equiv\sum_{i=1}^{K}p_{i}F_{i}(\boldsymbol{\theta}^{(t)})$ w.r.t. $\boldsymbol{p}$, when ignoring the constant and the response transformation is monotonically non-decreasing. In turn, along with the minimization objective in terms of parameter as in (1), we can see the overall objective of the federated system can be represented as the approximation of max-min optimization problem.
@@ 4.1 Improved Design for Better Decision Making
Additionally, when equipped with the negative logarithmic growth as a decision loss (i.e., eq. (4)), the problem c can be addressed due to its strict convexity and Lipscthiz continuity. (See Lemma 4.1) Note that when the loss function is convex, we can run the FTRL with a linearized loss (i.e., ~$\tilde{\ell}^{(t)}\left(\boldsymbol{p}\right)=\left\langle\boldsymbol{p},\boldsymbol{g}^{(t)}\right\rangle$ where $\boldsymbol{g}^{(t)}=\nabla\ell^{(t)}\left(\boldsymbol{p}^{(t)}\right)$). This is useful in that a closed-form update can be obtained thanks to the properties of the Fenchel conjugate (see Remark 4.5).
@@ 4.2.1 AAggFF-S: Algorithm for the Cross-Silo Federated Learning
In the cross-silo setting, it is typically assumed that all $K$ clients participate in $T$ rounds, since there are a moderately small number of clients in the federated system. Therefore, the server’s stateful decision making is beneficial for enhancing overall welfare across federation rounds. This is also favorable since existing OPS algorithms can be readily adopted.
@@ 4.2.2 AAggFF-D: Algorithm for the Cross-Device Federated Learning
The ONS has regret proportional to $K$ and runs in $\mathcal{O}\left(K^{2}+K^{3}\right)$ [^1] A generalized projection required for the Online Newton Step can be solved in $\tilde{\mathcal{O}}\left(K^{3}\right)$ (Agarwal et al., 2006; Hazan et al., 2007). per round, which is nearly impossible to be adopted for the cross-device FL setting due to large $K$, even though the logarithmic regret is guaranteed in $T$. Instead, we can exploit the variant of EG adapted to FTRL (Orabona, 2019), which can be run in $\mathcal{O}\left(K\right)$ time per round.
@@
The large number of clients coerces the federated system to introduce the client sampling scheme in each round. Therefore, the decision maker (i.e., the central server) cannot always observe all entries of a response vector per round. This is problematic in terms of OCO, since OCO algorithms assume that they can acquire intact response vector for every round $t\in[T]$. Instead, when the client sampling is introduced, the learner can only observe entries of sampled client indices in the round $t$, denoted as $S^{(t)}$.
@@
Denote $C=P\left(i\in S^{(t)}\right)$ as a client sampling probability in a cross-device FL setting for every round $t\in[T]$. The DR estimator $\breve{\boldsymbol{r}}^{(t)}$, of which element is defined in (10) is an unbiased estimator of given partially observed response vector $\boldsymbol{r}^{(t)}$. i.e., $\mathbb{E}\left[\breve{\boldsymbol{r}}^{(t)}\right]=\boldsymbol{r}^{(t)}$.
@@
Further denote $\breve{\boldsymbol{g}}$ as a gradient estimate from (11) using the DR estimator of a response vector according to (10), at an arbitrary reference $\boldsymbol{r}_{0}$. Then, $\breve{\boldsymbol{g}}$ is an unbiased estimator of the linearized gradient of a decision loss at $\boldsymbol{r}_{0}$, which is close to the true gradient: $\mathbb{E}\left[\breve{\boldsymbol{g}}\right]=\tilde{\boldsymbol{g}}\approx\boldsymbol{g}$.
@@ 6 Experimental Results
We conduct experiments on datasets mirroring realistic scenarios in federated systems: multiple modalities (vision, text, speech, and tabular form) and natural data partitioning. We briefly summarize FL settings of each dataset in Table 4. For the cross-silo setting, we used Berka tabular dataset (Berka, 1999), MQP clinical text dataset (McCreery et al., 2020), and ISIC oncological image dataset (Codella et al., 2018) (also a part of FLamby benchmark (Ogier du Terrail et al., 2022)). For the cross-device setting, we used CelebA vision dataset (Liu et al., 2015), Reddit text dataset (both are parts of LEAF benchmark (Caldas et al., 2019)) and SpeechCommands audio dataset (Warden, 2018).
@@ 7 Limitations and Future Works
Our work suggests interesting future directions for better federated systems, which may also be a limitation of the current work. First, we can exploit side information (e.g., parameters of local updates) to not preserve all clients’ mixing coefficients, and filter out malicious signals for robustness. For example, the former can be realized by adopting other decision making schemes such as contextual Bayesian optimization (Char et al., 2019), and the latter can be addressed by clustered FL (Ghosh et al., 2020; Sattler et al., 2020) for a group-wise estimation of mixing coefficients. Both directions are promising and may improve the practicality of federated systems. Furthermore, the FTRL objective can be replaced by the Follow-The-Perturbed-Leader (FTPL) (Kalai & Vempala, 2005), of which random perturbation in decision making process can be directly linked to the differential privacy (DP (Dwork, 2006)) guarantee (McMillan, 2018), which is frequently considered for the cross-silo setting. Last but not least, further convergence analysis is required w.r.t. the parameter perspective along with mixing coefficients, e.g., using bi-level optimization formulation.
@@ Acknowledgements
Federated Learning (FL) poses a potential risk of skewed performance distributions toward partial groups of clients. We believe that our proposed method can mitigate such an unfair situation in the federated system, thereby enhancing the trustworthiness and social welfare. A potential concern is the situation where malicious clients deliberately inflate their local signals to cause global updates to be biased towards themselves. To prevent such a catastrophic situation, we aim to improve the current work by exploiting information other than local losses, as mentioned in the Limitations and Future Works section. The authors are not aware of any other critical ethical/social implications otherwise, but are open to discussing them if they exist.
@@ C.1 Datasets
For the main experiments in Table 2 and Table 3, we used 3 datasets for the cross-silo setting (Berka (Berka, 1999), MQP (McCreery et al., 2020), and ISIC (Codella et al., 2018; Ogier du Terrail et al., 2022)), and 3 other datasets for the cross-device setting (CelebA, Reddit (Caldas et al., 2019), and SpeechCommands (Warden, 2018)). In this section, we describe details of each dataset about its task, metrics, and client partitioning. The statistics of all federated benchmarks are summarized in Table A2. Note that Avg. and Std. in the table refer to the average and a standard deviation of a sample size of each client in the federated system.
@@ 1. $\bullet$
    CelebA is a vision dataset containing the facial images of celebrities (Liu et al., 2015). It is curated for federated setting in LEAF benchmark (Caldas et al., 2019), and is targeted for the binary classification task (presence of smile). We follow the processing of (Caldas et al., 2019), thereby each client corresponds to each celebrity, having 9,343 total clients in the federated system. Finally, we used top-1 accuracy for the evaluation metric following (Caldas et al., 2019).
@@ D.1 Pseudocode for ClientUpdate
  Return: $F_{i}\left(\boldsymbol{\theta}\right)$, $\boldsymbol{\theta}-\boldsymbol{\theta}^{(E)}$.